Circular No. 9017

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 B3 (VAN NESS) G. Hug, Scranton, KS, U.S.A., reports his recovery of comet P/2005 R2

(cf. IAUC 8597) on images obtained with the 0.56-m reflector at his Sandlot Observatory. A very faint, straight, narrow tail was observed extending at least 5′–6′ in p.a. 300^{o}. The indicated correction to the prediction on MPC 62880 is ΔT = −0.06 day.

2009 UT α_{2000} δ_{2000} Mag. Observer
Jan. 31.36200 10^{h}57^{m}13.^{s}96 +0^{o}44′12′′.1 20.7 Hug

The above observation and three others, together with the following linked orbital elements and an ephemeris, appear on MPEC 2009-C03.

Epoch = 2011 June 8.0 TT

T = 2011 June 16.2308 TT  ω = 3.^{o}3321
e = 0.379597  Ω = 312.6719 }2000.0
q = 2.122533 AU  i = 10.2396

a = 3.421219 AU n^{o} = 0.1557517 P = 6.33 years

COMET P/2009 B4 (LINEAR) Hug also reports his recovery of P/2002 CW_{134} (cf. IAUC 7858) in

images taken at 25^{o} altitude; the comet did not appear different from stellar images in the same exposures. The indicated correction to prediction on MPC 56802 is ΔT = −0.32 day.

2009 UT α_{2000} δ_{2000} Mag. Observer
Jan. 31.48615 14^{h}30^{m}50.^{s}15 −22^{o}00′01′′.1 19.6 Hug
31.50481 14 30 51.92 −22 00 23.0 20.0  _{”}
Feb. 1.43053 14 32 21.53 −22 17 31.7 19.3  _{”}
1.45242 14 32 23.47 −22 17 56.4 19.5  _{”}
1.46701 14 32 24.91 −22 18 13.5 19.6  _{”}

The following linked orbital elements and an ephemeris appear on MPEC 2009-C04.

Epoch = 2009 Jan. 9.0 TT

T = 2009 Jan. 5.6142 TT  ω = 190.^{o}2682
e = 0.488719  Ω = 348.2590 }2000.0
q = 1.843732 AU  i = 15.2133

a = 3.606102 AU n^{o} = 0.1439286 P = 6.85 years

2009 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green